Problem 2 (1.1.5, 4 points): Find the rectangular parallelepiped of unit volume that has the minimum surface area. Hint: By eliminating one of the dimensions, show that the problem is equivalent to the minimization over > 0 and y > 0 of f(x,y) =xy +-+- Show that the sets {(x,y) I f(x,y) x > 0, y > 0} are compact for all scalars γ.
Answer #1

Assume a parallelopipe with dimensions: x,y,z

V=1= xyz

S= 2(xy+yz+zx)

z= 1/xy substitute z in S we get,

S= 2(xy+ 1/x+1/y)

argmin. S= argmin.{2(xy+ 1/x+1/y)}= argmin.(xy+1/x+1/z)

argmin. S= argmin.(\gamma (xy+ 1/x+1/y))= argmin.(xy+1/x+1/z)

for all \gamma as scaling doesnot change the value of the minimizing function